### v1

@@ Abstract
The category was introduced in [12] in order to provide a structural setting to the study of the many Gelfand-Tsetlin subalgebras appearing in the context of the diagrammatic Soergel category of Elias and Williamson [3]. The category have as objects, all that we called nil graded algebras associated to a triangular matrix and as morphisms, all the preserving degree homomorphisms of graded algebras between them. In this article we develop a series of isomorphism criteria on that we used later to find out relevant information on the Gelfand-Tsetlin subalgebras of the diagrammatic Soergel category.
@@ 1.1 Motivation
The diagrammatic Soergel Category was defined in 2016 by Elias and Williamson [3] in order to provide a diagrammatic approach to the study of the category of Soergel bimodules, a well known incarnation of the Hecke category, [18]. Since then, the category have gained attention of both algebraists and algebraic geometrist, due to its intrinsic connection with both branches of study.
@@
A particular interest is the study of the internal structure of the category itself, that is, the structure of the many Hom spaces and more particularly of the many End algebras. In this context, for example, the diagrammatic setting provided by Elias and Williamson have been crucial to find out and prove the cellularity of the Light leaves bases of Libedinsky [8],[3]. Another example of the power of this diagrammatic approach is the discovery due to Ryom-Hansen, of explicit Jucys-Murphy elements relative to the Libedinsky’s bases [19].
In [10], Libedinsky and Plaza, found out a hard connection between the category and what they called the Blob category, another diagrammatic category, inspired in the well known KLR-algebras of Khovanov, Lauda and Rouquier. The similar patterns that they observed in both categories motivate them to define the Blob vs Soergel conjecture, where roughly speaking, they declare that there is an explicit incarnation of the category (in type ), in terms of KLR-diagrams.
@@
Some of the first evidences on the veracity of the aforementioned conjecture can be found in the work of Espinoza and Plaza [4], where explicit isomorphisms were built between the corresponding Gelfand-Tsetlin subalgebras, that is, the algebras generated by the Jucys-Murphy elements, coming on one side from Two-color Soergel Calculus and on the other of certain idempotent truncation of the KLR version of the blob algebra (see [17]). In [13], Plaza, Ryom-Hansen and the author, obtain isomorphisms in a higher level, this time between certain endomorphism algebras of in type and the corresponding idempotent truncation of blob algebras. In both cases the procedure implies the search of optimal presentations in terms of generators and relations for the algebras involved, therefore, in both [4] and [13], the reader will find, not only a clear proof of each isomorphism, but also deep information on the structure of the corresponding algebras.
@@
Recently in [1], Bowman, Cox and Hazi have proved the Blob vs Soergel conjecture, they in fact, proved a more general result, obtaining as a consequence the equivalences predicted by Libedinsky and Plaza. The smart procedure used in [1], based in alcove geometry and paths configurations, allows the authors to cover all the cases in just one go, while in [4] and [13] only particular cases were covered. On the other hand the general approach developed in [1] do not implies the aforementioned internal information on the corresponding algebras obtained in [4] and [13] or any generalization of that. Therefore, we consider as an open problem the search of this internal information for all the cases no covered in [4] and [13], that is in principle, to find out optimal presentations for all the algebras related with the Blob vs Soergel conjecture.
@@
In the last few years, the author have been developing certain strategies to address this problem. For example in [11], presentations for the Gelfand-Tsetlin subalgebras of a family of idempotent truncations of the generalized blob algebras were found, later in [12], working on the side of the category the author developed a series of techniques and procedures to obtain optimal presentation for the Gelfand-Tsetlin subalgebras of the many End algebras for any type (far beyond the initial objective: type ). Both [11] and [12] are natural generalizations of [4] and particularly they seem to imply a hard connection between the light leaves bases of Libedinsky and the graded cellular bases built by Ryom-Hansen and the author in [14], just as was previously observed in [4] and [13], in that case, with respect to the graded cellular bases developed by Plaza and Ryom-Hansen in [17].
@@
The category was introduced in [12] in order to provide a structural setting to the study of the many Gelfand-Tsetlin subalgebras appearing in the context of the diagrammatic Soergel category The category have as objects, all that we called nil graded algebras associated to a triangular matrix and as morphisms, all the preserving degree homomorphisms of graded algebras between them. Naturally, the aforementioned Gelfand-Tsetlin subalgebras of are (the main) objects in our category, and all the knowledge of developed in [12] allowed us to find out the desired presentations for those algebras. On the other hand, the category is by itself an interesting object of study and the results obtained in [12] open the door to a deeper analysis on it.
Being more specific, the objects of the category are commutative graded algebras defined by a presentation with a finite set of generators and certain quadratic relations codified by a strictly triangular matrix. In [12] we define and make a first study of those algebras, obtaining a series of techniques and corresponding applications to the study of the many Gelfand-Tsetlin subalgebras of the category Since each objects of is totally determined by a codifying triangular matrix, it seem plausible to ask when two of those objects are isomorphic and which role play the corresponding codifying matrices in this aspect. To answer this question is the main motivation of this and other works we have in progress.
In section 2, we develop a series of isomorphism criteria on that partially answer the above question. In Theorem 2.2 we obtain an absolute (but not easy to apply) matrix criterion to determine if two objects are or not isomorphic, this is the key result that allowed us to find out different approaches to a complete classification of the objects of Applying conveniently theorem 2.2, we could describe completely the class of the algebra associated to the zero matrix obtaining a very clear and easy to apply criterium (Theorem 2.9) to determine if the algebra is isomorphic to or not, depending on the entries of the codifying matrix The case of the matrix although very particular, can be cleverly used to determine if other two objects of are or not isomorphic (see for example the proof of proposition 3.3). Another fundamental result of this article is Theorem 2.39, where we describe another absolute criterium (again, not always easy to apply) to determine if two objects are or not isomorphic. In the case of theorem 2.39, the criterion states that two objects and are isomorphic if and only if, there is a finite sequence of certain matrix operations (that we called elementary triangular operations), that progressively transform the matrix into the matrix
Our first examples are developed in subsection 2.2, where the smaller cases were studied. The results that we found there inspire all the analysis we did in this article. In particular, they open interesting questions for further works, for example, since we observe that for small matrices, we have a very small number of isomorphic classes in then is natural to ask:
@@
Although we do not have an answer for this question yet, in subsection 2.4, we obtain a lower bound for the number of classes depending on the size of the codifying matrices. In theorem 2.23, we describe an explicit set of representatives of different classes, for each size We left Q1 open, together with the search of a complete set of representatives and explicit criteria to describe any isomorphic class such as we did in the case of the matrix
On the other hand, in section 3, we advocate our attention to apply the results developed in section 2, to the study of the Gelfand-Tsetlin subalgebras coming from the diagrammatic Soergel category In [12], for each element we a define the SLTM, and the corresponding nil graded algebra We have proved in [12], that the algebra corresponds to the Gelfand-Tsetlin subalgebra of the endomorphism algebra in the diagrammatic Soergel Category of Elias and Williamson, relative to the Libedynski’s light leaves basis and the set of Jucys-Murphy elements defined by Ryom-Hansen (See [3],[8] and [19]).
In [12] we show that for the case of a Coxeter system of type the algebras are invariant under commuting moves. That result naturally open the following two questions:
@@
Is the algebra invariant under commuting moves in any other type?
- Q3:
Is the algebra invariant under other braid moves?
In section 3, we answer those questions applying the different techniques developed in section 2. In theorem 3.2, considering the geometric realization of a Coxeter system (assuming all the technical conditions described in [3]), we prove that the algebra is invariant under commuting moves, independently of the particular type of the system. Although we already prove this for type in [12], the proof is different in the general case. In theorem 3.2 we use the new techniques developed in this article. Finally in proposition 3.3 we answer We provide an explicit example and again, we use the techniques developed in section 2 to verify that the algebra is not always invariant under other braid moves.
1.2 Acknowledgment
This work was supported by FONDECYT de Postdoctorado 2024, N°3240046. ANID, Chile.
@@ 1.3 Notations and terminology
is a fixed field of characteristic different from
is the zero matrix.
@@
Given a SLTM, a SLTM, and a matrix we define the matrix by blocks, as follows:
@@ 1.4 Generalities
2 Isomorphism criteria
2.1 The fundamental equations
We shall assume that is a SLTM, is a SLTM, and are the generators defining and respectively. A preserving degree homomorphism defines a matrix by the equation
The following lemma provides a first matrix criterium to find out if a given matrix defines a homomorphism or not.
@@ 2.1 The fundamental equations
Again by relation 1.4.1, we obtain
@@
for any and any pair such that as desired. ∎
The following theorem is a direct consequence of lemma 2.1. In this case we shall assume that and are both SLTM.
@@
Given a matrix, then the assignment
defines a preserving degree isomorphisms of algebras if and only if
and is an invertible matrix.

### v2

@@ Abstract
The category was introduced in [15] in order to provide a structural setting to the study of certain commutative algebras appearing in the context of the diagrammatic Soergel category of Elias and Williamson [3]. The category have as objects, all that we called nil graded algebras associated to triangular matrices and as morphisms, all the preserving degree homomorphisms of graded algebras between them. In this article we develop a series of isomorphism criteria on and we show some of the applications relative to the search of a complete classification.
keywords: Commutative graded algebras, Jucys-Murphy elements, Soergel Calculus.
@@ 1.1 Motivation
The diagrammatic Soergel Category is a linear and monoidal category, it was defined in 2016 by Elias and Williamson [3] in order to provide a diagrammatic approach to the study of the category of Soergel bimodules, a well-known incarnation of the Hecke category, [21]. Since then, the category have gained attention of both algebraists and algebraic geometrist, due to its intrinsic connection with both branches of study. For further studies on this topic, we recommend to see [4] and [12] for example.
@@
A particular interest is the study of the internal structure of the category itself, that is, the structure of the many Hom spaces and more particularly of the many End algebras. In this context, for example, the diagrammatic setting provided by Elias and Williamson have been crucial to find out and prove the cellularity of the Light leaves bases of Libedinsky [9],[3]. Another example of the power of this diagrammatic approach is the discovery due to Ryom-Hansen, of explicit Jucys-Murphy elements relative to the Libedinskyβs bases [22] (In the following when we mention Jucys-Murphy elements in we refer to those developed by Ryom-Hansen in [22]). Finally we would like to mention, probably the most important application of this diagrammatic setting, the disproof of the famous Lusztigβs conjecture, given by Williamson in [24].
@@
In [13], Libedinsky and Plaza, found out a connection between the category and what they called the Blob category, another diagrammatic category, inspired in the well-known KLR-algebras of Khovanov, Lauda and Rouquier. The similar patterns that they observed in both categories motivate them to define the Blob vs Soergel conjecture, where roughly speaking, they declare that these two categories are equivalent. Some of the first evidences on the veracity of the aforementioned conjecture can be found in the work of Espinoza and Plaza [5], where explicit isomorphisms were built between the subalgebras generated by the Jucys-Murphy elements, coming on one side from Two-color Soergel Calculus and on the other of certain idempotent truncation of the KLR version of the blob algebra (see [20]). In [16], Plaza, Ryom-Hansen and the author, obtain isomorphisms in a higher level, this time between endomorphism algebras of in type and the corresponding idempotent truncation of blob algebras. In both cases the procedure implies the search of adequate presentations in terms of generators and relations for the algebras involved, therefore, in both [5] and [16], the reader will find, not only a clear proof of each isomorphism, but also more information on the structure of the corresponding algebras.
@@
Some years ago, Bowman, Cox and Hazi have proved in [1], the Blob vs Soergel conjecture, they in fact, proved a more general result, obtaining as a consequence the equivalences predicted by Libedinsky and Plaza. The procedure used in [1], based in alcove geometry and paths configurations, allows the authors to cover all the cases in just one go, while in [5] and [16] only particular cases were covered. On the other hand the general approach developed in [1] do not implies the aforementioned internal information on the corresponding algebras obtained in [5] and [16] or any generalization of that. Therefore, we consider as an open problem the search of this internal information for all the cases no covered in [5] and [16], that is, in principle, to find out presentations for all the algebras related with the Blob vs Soergel conjecture.
@@
In the last few years, we have been developing certain strategies to address this problem. For example in [14], we obtained presentations for the subalgebras generated by Jucys-Murphy elements in certain family of idempotent truncations of the generalized blob algebras. Later in [15], working on the side of the category we developed a series of techniques and procedures to obtain presentations for the subalgebras generated by Jucys-Murphy elements (from now the algebras), in the many End algebras for any type (far beyond the initial objective: type ). Both [14] and [15] are natural generalizations of [5].
In general the subalgebra generated by the Jucys-Murphy elements in a cellular algebra is a commutative subalgebra that under certain conditions, it could be a maximal commutative subalgebra of (see [18]). One can expect that subalgebra contains important information on the algebra itself. On the other hand Jucys-Murphy elements are very relevant in the study of the Representation Theory of the corresponding cellular algebra (see [18]). We think that all the study that we are doing on algebras in the context of the category will help to learn more on this important category.
@@
The category was introduced in [15] in order to provide a structural setting to the study of the many algebras appearing in the context of the diagrammatic Soergel category The category have as objects, all that we called nil graded algebras associated to triangular matrices and as morphisms, all the preserving degree homomorphisms of graded algebras between them. Naturally, the aforementioned algebras of are (the main) objects in our category, and all the knowledge of developed in [15] allowed us to find out the desired presentations for those algebras. In particular we would like to use our work in the search of a complete classification of the many algebras of the Soergel category
On the other hand, the category is, in our opinion, by itself an interesting object of study, since their objects are basically certain quotients of polynomial algebras on several variables, they have potential connections with other branches of study, for example with Algebraic Geometry. To obtain tools to classify the objects on could help, for example, to obtain new tools for analysis of certain types of Algebraic Varieties.
Being more specific, the objects of the category are commutative graded algebras defined by a presentation with a finite set of generators and certain quadratic relations codified by a strictly lower triangular matrix. The dependence on the codifying matrices, invite us to ask how those matrices, could be relevant when we want to know if two of those algebras are isomorphic or not. Moreover, we would like to know if a complete classification of the objects of is possible, and which role play the corresponding codifying matrices in these aspects. In this article we give some first steps in the search of answers for those questions. The content of this article is organized as follows:
- β’
β’
In subsection 2.2, we obtain our first isomorphism criterion. In Theorem 2.6 we provide an absolute (but not easy to apply) matrix criterion to determine if two algebras in are or not isomorphic. Theorem 2.6, basically asserts that two algebras and in the category are isomorphic if and only if there exist an invertible matrix whose entries satisfy certain system of quadratic equations, determined by the entries of and
β’
In subsection 2.3, we define a set of matrix operations, that we called Elementary triangular operations (or ETOβs for shortness). This matrix operations act on the set of triangular matrices, each one subject to certain restrictions. In Definition 2.17, we define Equivalence by ETOβs, an equivalence relation between triangular matrices. In Theorem 2.19, we show that if two matrices are equivalent by ETOβs, then the corresponding algebras are isomorphic as objects of the category
β’
In Proposition 2.20, we connect the matrix coming from Theorem 2.6 and ETOβs. We prove that if is any elementary matrix, defining an isomorphism then it implies that the matrices and are equivalent by ETOβs. Since any isomorphism is determined by an invertible matrix (Theorem 2.6), and any invertible matrix can be factored as a product of elementary matrices, it seem sensible to think that the existence of an isomorphism between and will imply that the matrices are necessarily equivalent by ETOβs. We declare this in Conjecture 2.21. We do not have a proof of this yet, but we have an important set of evidences on its veracity, we show some of them in this article.
β’
In Section 3 we show some applications of the criteria obtained in Section 2. In Subsection 3.1, we obtain a complete classification of algebras where is a lower matrix of size or (see Propositions 3.1 and 3.3). We also prove for those cases the veracity of Conjecture 2.21 (see Corollaries 3.2 and 3.4).
β’
In Subsection 3.2, we use Theorem 2.6 to deduce an absolute criterion, depending explicitly on the entries of a matrix to determine if the algebras and are isomorphic or not. Theorem 3.12, is a clear example of what we were talking about, when we propose ourselves to look for criteria coming from the codifying matrices, to determine when two algebras in are isomorphic or not. Although it only cover one isomorphism class, we consider this result as a challenging invitation to obtain such as explicit criteria for other isomorphism classes. This is the motivation to keep on working in the search of a complete classification of the objects in On the other hand Proposition 3.14 provide another evidence on the veracity of Conjecture 2.21. In this case, Proposition 3.14 asserts that if an algebra is isomorphic to then necessarily the matrices and are equivalent by ETOβs.
β’
As we mentioned above, in Subsection 3.1, we obtained a complete classification for algebras codified by matrices of size or In fact, Propositions 3.1 and 3.3, provide explicitly the number of classes for those sizes. In both cases, the number is finite and it is independent on the ground field. Those facts invite us to open the following questions:
@@
Is the number always independent of the ground field?
We do not have answer for those questions yet, but we expect to address this problem in future works. For instance, motivated uniquely by our intuition, we define Conjecture 3.17, where we state that the number of isomorphism classes for depends on the ground field. Moreover the number is finite if and only if the ground field is finite.
β’
Independently on if the number is finite or not, in Subsection 3.3, we obtain a lower bound for the number of classes for each size that is independent on the ground field. In Theorem 3.16, we describe an explicit set of representatives of different classes. This result will be useful, when we have answered Q1 and Q2, and naturally look for the exact value of
The main results of this article are:
@@ 1.2 Notations and terminology
The following notation will be used along this article:
β’
is a fixed field of characteristic different from We also denote
β’
is the zero matrix and is the identity matrix, both of size
β’
@@
β’
Given a SLTM, an scalar and indices we denote by the element When there is no possible confusion we only write instead of
1.3 Acknowledgement
It is a pleasure to thank Steen Ryom-Hansen for stimulating conversations on Jucys-Murphy elements and algebras. I would specially want to thank the anonymous referee for communicating detailed suggestions and remarks that helped to improve the clarity and presentation of this article.
This work was supported by:
β’
FONDECYT de Postdoctorado 2024, NΒ°3240046, ANID, Chile.
β’
SubvenciΓ³n a la InstalaciΓ³n en la Academia 2024, NΒ°85240053, ANID, Chile.
2 Isomorphism criteria
@@ 2.1 Generalities
2.2 Fundamental equation
We shall assume that is a SLTM, is a SLTM, and are the generators defining and respectively.
Note that any morphism defines a matrix by the equation
In particular any morphism is totally determined by a linear transformation On the other hand, a given linear transformation not necessarily defines a morphism in the category The following lemma provides a first matrix criterion to find out if a given matrix defines a morphism via equation 2.2.1, or not:
@@ 2.2 Fundamental equation
Again by relation 2.1.1, we obtain
@@
for any and any pair such that as desired. β
@@
Let be two SLTM. Given a matrix, then the assignment
defines an isomorphisms if and only if is an invertible matrix whose entries satisfy the following system of equations:
Proof.
If is an isomorphism, then clearly the associated linear transformation (obtained by restriction) is invertible, and therefore the matrix (defined by equation 2.2.1) is invertible. Lemma 2.5 implies that the entries of satisfy equation 2.2.6.
On the other hand if is an invertible matrix, and satisfies equation 2.2.6, then by Lemma 2.5, it defines a morphism Let be the inverse matrix of Let us define the function as follows
where the sum runs over all the with is any scalar and is given by the equation
Is not difficult to see that is the inverse function of Since is in particular an homomorphism of algebras, we conclude directly that is also an homomorphism of algebras. Since preserves degrees, so does therefore is a morphism in the category and this implies that is an isomorphism in the category β
2.3 Elementary triangular operations
In this section we define three matrix operations, that we call elementary triangular operations (ETOβs). Basically each ETO transform a SLTM, , into a new SLTM, As we will see in Definitions 2.7, 2.10 and 2.13, each ETO has certain restrictions, that is, we only are allow to apply them on matrices that satisfies a series of conditions. Anyway, we shall prove that, whenever this transformation is well-defined, we obtain a matrix such that Moreover we shall prove that if there is a finite sequence of, well-defined, ETOβs that transform into then we have that
Definition 2.7.
Let be a SLTM, fix an and an scalar We define the matrix as the SLTM, whose entries are given by:
Example 2.8.
Consider the matrix
then
Lemma 2.9.
Let be a SLTM, fix and an scalar Then
Proof.
Let (as in definition 2.7). Let and the generators of and respectively. We claim that the assignment
defines an isomorphism To see that, we only have to check that
for each and that the associated matrix is invertible. Now note that
- 1.
- 2.
- 3.
By the last analysis, we conclude that the assignment defines an homomorphism Is not difficult to see that the associated matrix is invertible, in fact, the entries of are given by
that is, then and we conclude that defines an isomorphism β
In the following, we shall denote by the associated matrix of the isomorphism defined in equation 2.3.1. For example if and we have:
Note that is an elementary matrix.
Definition 2.10.
1.
The entry for each
2.
The entry for each
then we define the matrix as the SLTM, whose entries are given by:
Example 2.11.
Consider the matrix
and take and Is not difficult to see that satisfies the conditions of Definition 2.10. Then the matrix is well-defined. Moreover we have
Lemma 2.12.
Let be a SLTM, let If satisfies the conditions of Definition 2.10, then
Proof.
Let (as in definition 2.10). Let and the generators of and respectively. We claim that the assignment
defines an isomorphism To see that, we only have to check that
for each and that the associated matrix is invertible. Now note that
- 1.
- 2.
- 3.
4.
For we have on one side
- 5.
The last analysis implies that the assignment defines a homomorphism Is not difficult to see that the associated matrix is invertible, in fact, the entries of are given by
Therefore and we conclude that defines an isomorphism β
In the following we shall denote by the associated matrix of the isomorphism defined in equation 2.3.7. For example if and we have:
Note that is an elementary matrix.
Recall that for a given SLTM, we denote (see subsection 1.2).
Definition 2.13.
1.
2.
and for each
Then we define the matrix as the SLTM, whose engtries are given by:
Example 2.14.
Consider the matrix
Example 2.15.
Therefore we have
Lemma 2.16.
Proof.
We shall assume that The case is easier and it is left to the readers.
Assuming that and for each then the matrix is well-defined and its entries are given by equation 2.3.11 (see Definition 2.13). Let and the generators of and respectively. We claim that the assignment:
defines an isomorphism As in Lemmas 2.9 and 2.12, we only have to check that
for each and that the associated matrix is invertible. Now note that:
- 1.
2.
If we have on one side:
The last step is because equation 2.3.11, from where we know that for all for and Also by the restrictions given in definition 2.13, we have that for each
On the other side, equation 2.3.12 implies that
Note that by our hypothesis, we have then we can reduce the expression on the parenthesis in the first sum to obtaining:
then equation 2.3.13 holds in this case.
3.
If we have on one side:
since by equation 2.3.11, that for and
By the above analysis we conclude that, the assignment defines an homomorphism Is not difficult to see that the matrix is invertible, in fact, the entries of are given by:
therefore and we conclude that defines an isomorphism. β
In the following we shall denote by the associated matrix of the isomorphism defined in equation 2.3.7. For example if and we have:
Note that is an elementary matrix.
Definition 2.17.
We call Elementary triangular operations (ETO for short) any of the following matrix assignments (whenever they are well-defined):
Given two SLTM, and we say that they are equivalent by ETO if there is a finite sequence of ETOβs that transform (progressively) into We denote by whenever and are equivalent by ETO.
It is clear that defines an equivalence relation on the set of all SLTM. In particular note that
Each of the assignments in equation 2.3.17 is well-defined if and only if the corresponding assignment in equation 2.3.16 is well-defined.
Example 2.18.
Consider the matrices
then In fact one can check that
Theorem 2.19.
Let be two SLTM. If then
Proof.
If we assume that that is, there is a finite sequence of ETOβs, that transform into say where each is one of the three ETOβs described in Definition 2.17, and the corresponding matrix is well-defined, that is, respects the corresponding restrictions of the operation Then Lemmas 2.9, 2.12 and 2.16 imply that at each step The result follows by transitivity of relation
β
Proposition 2.20.
Let be two SLTM. Assuming that there is an isomorphism such that, the associated matrix is an elementary matrix, then
Proof.
We fix a SLTM, T, and we left the matrix variable. Let be an elementary matrix. We separate our analysis in the following cases:
- Case 1
If for some and some (for notation, see just below of the proof of Lemma 2.9). If we assume that defines an isomorphism then the entries of must satisfy equation 2.2.6, where are replaced by the entries of our fixed matrix and are replaced by the entries of which are given explicitly in equation 2.3.4. Using equation 2.2.6 in this way, we have:
- 1.
- 2.
3.
For any other replacement of a triple such that we obtain equations implying that or equations of the form
The above analysis implies that the entries of are given by equation 2.3.1, and then Therefore in this case.
- Case 2
If for some (for notation, see just below of the proof of Lemma 2.12). We apply a similar analysis as in Case 1, we assume that defines an isomorphism then the entries of must satisfy equation 2.2.6, where are replaced by the entries of our fixed matrix and are replaced by the entries of which are given explicitly in equation 2.3.9. Using equation 2.2.6 in this way, we have:
- 1.
- 2.
- 3.
- 4.
- 5.
- 6.
7.
If we replace a triple such that and in equation 2.2.6, we obtain
(2.3.26) -
8.
If we replace a triple such that and in equation 2.2.6, we obtain
(2.3.27) -
9.
If we replace a triple such that and and in equation 2.2.6, we obtain
(2.3.28) -
10.
If we replace a triple such that and and in equation 2.2.6, we obtain
(2.3.29) -
11.
For any other replacement of a triple such that we obtain equations implying that or equations of the form
Note that equations 2.3.20, 2.3.21 and 2.3.22 imply that if the matrix defines an isomorphism then necessarily, the matrix must satisfy all the restrictions described in Definition 2.10. The above analysis imply that and therefore
- Case 3
If (for notation see just below of the proof of Lemma 2.16). The analysis is similar to those of Case 1 and Case 2, we have to replace the entries (defined in equation 2.3.14), in equation 2.2.6, collecting restrictions on and defining equations for the entries of We shall assume that the case is easier and it is left to the readers.
In this case we have the following situations:
- 1.
2.
If we replace a triple such that and in equation 2.2.6, we obtain
(2.3.31) - 3.
- 4.
5.
If we replace a triple such that in equation 2.2.6, we obtain
(2.3.35) -
6.
For any other replacement of a triple such that we obtain equations implying that or equations of the form
Note that equations 2.3.30, and 2.3.33 imply that if the matrix defines an isomorphism then necessarily, the matrix must satisfy all the restrictions described in Definition 2.10. The above analysis imply that and therefore
- Case 4
If then necessarily that is, there is a pair and an scalar such that
(2.3.36) Working exactly as we did in the three previous cases, we obtain that:
1.
The entries of have the following restrictions:
(2.3.37) In particular, they imply that for all and therefore the matrix is well-defined.
2.
The entries of depends on as follows:
Now note that the entries of the matrix are given by:
(2.3.39) In particular we have that for all and for all This implies that the matrix is well-defined.
Now note that the entries of are given by:
(2.3.40) Let us define the matrix Since the operation have no restrictions, is well-defined and its entries are given by:
(2.3.41) Note in particular that for all therefore we can define the matrix whose entries are given by:
(2.3.42) Finally we define the matrix By contruction and one can check that the entries of are given by
(2.3.43) Comparing equations 2.3.39 and 2.3.43, one conclude that and therefore as desired.
1.
β
Note that in Case 4 of the proof of Proposition 2.20, the matrix can be factored as follows:
and that factorization yields the following well-defined process of transformations:
That is, the restrictions and relations for matrices obtained, when the entries of are replaced in the system of equations 2.2.6, are coherent with the process of transformations induced by the factorization given in equation 2.3.44.
More generally, any invertible matrix can be factored as a product of elementary matrices, and each elementary matrix can be written as product of matrices of the forms or Now if we assume that is an invertible matrix, that define an isomorphism for certain variable matrices then the system of equations 2.2.6, will imply certain restrictions and relations on and Based in the proof of Proposition 2.20, it is sensible to think that, there will be a factorization with matrices of one of the forms or yielding a well-defined process of transformations:
At this point we do not have a well proof of that fact, but in an enormous set of explicit examples, we always obtain evidences of its veracity. Based on that, we define the following:
Conjecture 2.21.
(Isomorphism vs Equivalence by ETOβs) Let be two SLTM. If then
3 Applications
In this section we apply the isomorphism criteria obtained in Section 2, to give some first steps in the search of a complete classification of the objects of the category